Sunday, June 20, 2010

Horseshoe Fire Briefing & Map

At this morning's briefing in Portal, fire team leaders summarized yesterday's work and laid out the tasks for today. The fire's behavior yesterday was somewhat unexpected, because the big cliffs failed to delay it as much as had been anticipated. As the new fire map shows, the fire continues to burn north along the ridge in the high Chiricahua Mountains, southwest of Portal Peak. Over 3000 acres have burned, at a cost of nearly $9 million. The fire is actively burning over roughly 400 acres and remains 25% contained.

Blue arrow shows where water bars will be put into place today to moderate erosion along the fire control line.

In about 2 or 3 days, the team may move to moderate the downslope spread of the fire into lower South Fork, but at the moment that action is premature. This map shows a number of MAP, or Management Action Point, lines. When the fire crosses those lines, it triggers various responses from the fire team. Many of the MAP lines shown here are new and have been designated since the Type 1 team gave us a summary of the initial seven lines: I'll have to find out what the new lines represent. 

The MAPs help to guide strategy: if the fire were to reach MAP 3, it could have the future consequences of possible loss of structures and unacceptable resource damage, and would likely trigger evacuation notices for cabins and homes from South Fork to Portal. Obviously no one wants that to happen, and I'm guessing that new containment work being planned would be designed to prevent the fire's spread to that and other MAP lines. The fire's reaching MAP 5 would result in unacceptable damage to the South Fork watershed, and the fire team's work above Maple Camp (where the fire is now contained) has thus far prevented that from happening.

Included in this fire map are activity centers for several pairs of Spotted Owls; special care is being taken there to moderate the effects of the fire.

The crews have finished most of the water bars in Divisions A and B of the fire perimeter, and today will start building water bars in Division R (shown on the map). Chipping operations along the main Cave Creek Road are finished––and looking very good!––and cut firewood is being hauled to the Visitor Center, where the rumor is that it will be sold. Brushing and chipping continue along the South Fork road and picnic area.

The South Fork and Herb Martyr roads remain closed.

Today's weather conditions could spawn a dry thunderstorm, and crews were cautioned that lightning is a hazard.

Yesterday a fire team rescued a lost boy who had become separated from his parents when he went in search of ladybugs (obviously a lad whose passion is right in synch with that of many PortaleƱos!).

Woodpeckers in the Chiricahuas:
Acorn, Ladder-backed, Hairy & Arizona
(Watercolor & Gouache by Narca)

No comments:

Post a Comment